# HARP: Hallucination Detection via Reasoning Subspace Projection

### v1

@@ 1 Introduction
Large Language Models (LLMs) have recently demonstrated remarkable generative capabilities and broad applicability across various natural language processing tasks [1, 2, 3]. However, hallucinations—i.e., model-generated information inconsistent with objective facts—remain a major obstacle to their deployment in critical decision-making scenarios [4, 5]. Consequently, efficiently and accurately detecting hallucinations during LLMs generation has become a pressing challenge.
From a cognitive perspective, the hallucination behavior of LLMs is to some extent similar to human’s “nonsense” behavior. When answering complex questions, humans typically follow a “Reasoning $\rightarrow$ Expression” process: they first perform internal reasoning and then express part of the thought outcomes in language [6]. Therefore, although assessing the veracity of the answer is challenging when based solely on linguistic output, it can be substantially improved by observing the complete reasoning process [7]. By analogy, achieving high-precision hallucination detection in LLMs requires placing greater emphasis on the reasoning information encoded within the hidden states, rather than primarily on the semantic content of the outputs.
@@ 2 Related Work
Mechanistic interpretability of LLMs. Research on mechanistic interpretability mainly focuses on two aspects: model parameters and hidden states. For the former, several works analyze weight matrices to uncover structural properties and interactions among modules. For instance, Merullo et al.[8] and Cheng et al.[9] employ SVD to characterize attention-head structures and investigate their roles in downstream tasks. However, such approaches remain largely at the structural level, offering limited semantic interpretability. To address this, recent work has shifted toward analyzing hidden states directly, probing the predictive relationship between intermediate representations and model outputs[10, 11, 12, 13, 14].
Hallucination detection. The success of probing methods has motivated researchers to adopt similar ideas in hallucination detection[15, 16, 17]. For instance, Du et al.[18] leverage unlabeled embeddings and apply SVD to identify key subspace directions, followed by probing to link these directions to hallucinations. Yet, probing-based methods often rely on predefined supervised labels, making them less generalizable when feature dimensions are large or category priors are incomplete. Another line of work approaches hallucination detection from the perspective of output consistency. Chen et al.[19] quantify semantic agreement across multiple generations via covariance eigenvalues, while Farquhar et al.[20] utilize clustering and semantic entropy to detect hallucinations. These methods are effective in practice but may suffer from misclassification due to their inability to exploit internal reasoning information.
@@ 4.3 Construction of Semantic and Reasoning Subspaces via Low-Rank Approximation
Specifically, based on Equation 7, for any $k<rank(W_{unemb})$, the Eckart–Young–Mirsky theorem [21, 22] gives the best rank-$k$ approximation $W_{k}$ of $W_{unemb}$ under the Frobenius norm as:
@@ 4.4 Hallucination Detection via Reasoning subspace Projection
where $\theta$ denotes the parameters of the hallucination detector. ${{g}_{\theta}}\left({proj}_{R}\left(h_{l}^{(i)}\right)\right)$ represents the hallucination score of the $i$-th token, and $g_{\theta}\left(y|x\right)\in\left[0,1\right]$ is the score for the entire QA pair. We optimize the detector using the Binary Cross-Entropy Loss [23]:
@@ 5.1 Experimental Setup
Datasets and models. Our experiments cover four generative question answering (QA) tasks, including three open-domain dialogue QA datasets—NQ Open[24], TruthfulQA[25] (generation task), and TriviaQA[26]—and one reading comprehension dataset, TyDiQA-GP (English)[27]. To assess the effectiveness and generality of our proposed framework, we conduct evaluations using two widely adopted open-source foundation models: Qwen-2.5-7B-Instruct[1] and LLaMA-3.1-8B[2]. More dataset and inference details are provided in Appendix A.
@@
Baseline Methods. HARP is compared with several mainstream hallucination detection methods, including Perplexity[28], LN-Entropy[29], Semantic Entropy[20], Lexical Similarity[30], EigenScore[19], and HaloScope[18].
Correctness Measurement. Following Chen et al.[19], correctness is determined based on ROUGE-L and semantic similarity between generated and reference answers. Semantic similarity is computed using the BLEURT model[31, 17]. An answer is considered correct if its ROUGE-L score exceeds 0.7 or its semantic similarity exceeds 0.5.
@@
[…]
@@
[…]
@@ 5.2 Main Results
Table 1 summarizes the AUROC scores (in %) of various hallucination detection methods across four QA datasets, using Qwen-2.5-7B-Instruct and LLaMA-3.1-8B as backbone models. Several key findings emerge from these results. (1) HARP consistently outperforms all baseline methods across all datasets and models, often by a significant margin. For instance, on TriviaQA, HARP achieves AUROC scores of 92.8% on Qwen and 92.9% on LLaMA, yielding improvements of +7.5% and +16.6%, respectively, over the second-best method, demonstrating its robustness and scalability across architectures and data characteristics. (2) Baseline methods such as Perplexity and HaloScope perform competitively on simpler datasets like TriviaQA, where answers are often limited to one or two tokens, but their performance deteriorates sharply on more complex datasets such as TyDiQA, which contains long contexts and accompanying documents. In contrast, HARP maintains high AUROC scores of 88.4% on Qwen and 86.6% on LLaMA in these challenging settings, highlighting its ability to handle reasoning-intensive and context-rich inputs. (3) Sampling-based methods, such as Semantic Entropy, Lexical Similarity, and EigenScore, incur higher computational costs but still fail to achieve comparable performance, whereas HARP’s single-pass approach provides both superior efficiency and accuracy. Collectively, these findings validate the effectiveness, robustness, and practical utility of HARP for hallucination detection in diverse QA scenarios.
@@ 5.3 More Analysis
Selection of Hallucination Score Threshold. In practice, it is necessary to set a hallucination score threshold $\alpha$ so that $\widehat{G}(y|x)$ produces a clear binary decision. As shown in 6(a), when $\alpha$ is between 0.2 and 0.8, detection accuracy remains high, indicating a substantial separation between normal and hallucinated answers under $\widehat{G}$. To align with common expectations for a binary classifier, we set $\alpha=0.5$, where $\widehat{G}(y|x)=\mathbb{I}\left[g_{\theta}(y|x)>0.5\right]$.
Figure 6: (a) Effect of hallucination score threshold on detection accuracy. (b) Cross-dataset generalization. “(s)” indicates the source dataset used for training the hallucination detector; “(t)” indicates the target dataset.
Robustness Experiments. To apply HARP in real-world scenarios, we examined its performance under distribution shifts between training and test sets. We trained the hallucination detector on a source dataset $s$ and evaluated it on different target datasets $t$. 6(b) shows that HARP generalizes well across multiple target datasets. Notably, when trained on TriviaQA, its accuracy on NQ Open is nearly identical to directly training on NQ Open, demonstrating HARP’s strong robustness and cross-distribution adaptability.
@@ 6 Conclusion
In this study, we introduced HARP, a novel hallucination detection method that leverages only reasoning information as input features, achieving high detection accuracy while maintaining strong robustness. First, we showed that the hidden state space admits a direct-sum decomposition into a semantic subspace and a reasoning subspace, and that the Unembedding layer can effectively separate these two components. Building on this, we applied singular value decomposition to the parameters of the Unembedding layer and, following the Eckart–Young–Mirsky theorem, approximated $W_{unemb}$ with its best rank-$k$ approximation $W_{k}$. Setting $k=d\times 95\%$, we identified basis vectors for both the semantic and reasoning subspaces that align with empirical observations. Finally, HARP constructs an accurate and efficient hallucination detector by using the projections of hidden states in the reasoning subspace as input features. Experiments show that HARP significantly outperforms existing mainstream hallucination detection methods and maintains robustness under distribution shifts across datasets. In addition, we present a proof-of-concept demonstration of hallucination mitigation using our framework in Appendix D and aim to inspire future research in this direction.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have recently demonstrated remarkable generative capabilities and broad applicability across various natural language processing tasks Qwen2.5 ; Llama_3 ; LLM_Survey . However, hallucinations—i.e., model-generated information inconsistent with objective facts—remain a major obstacle to their deployment in critical decision-making scenarios ji2023survey ; Huang2025Survey . Consequently, efficiently and accurately detecting hallucinations during LLMs generation has become a pressing challenge.
From a cognitive perspective, the hallucination behavior of LLMs is to some extent similar to human’s “nonsense” behavior. When answering complex questions, humans typically follow a “Reasoning $\rightarrow$ Expression” process: they first perform internal reasoning and then express part of the thought outcomes in language Mental_models . Therefore, although assessing the veracity of the answer is challenging when based solely on linguistic output, it can be substantially improved by observing the complete reasoning process frank2012predicting . By analogy, achieving high-precision hallucination detection in LLMs requires placing greater emphasis on the reasoning information encoded within the hidden states, rather than primarily on the semantic content of the outputs.
@@ 2 Related Work
Mechanistic interpretability of LLMs. Research on mechanistic interpretability mainly focuses on two aspects: model parameters and hidden states. For the former, several works analyze weight matrices to uncover structural properties and interactions among modules. For instance, Merullo et al.merullo2024talking and Cheng et al.cheng2024enhancing employ SVD to characterize attention-head structures and investigate their roles in downstream tasks. However, such approaches remain largely at the structural level, offering limited semantic interpretability. To address this, recent work has shifted toward analyzing hidden states directly, probing the predictive relationship between intermediate representations and model outputsgurnee2023finding ; lv2024interpreting ; ju2024large ; he2024decoding ; jin2025exploring .
Hallucination detection. The success of probing methods has motivated researchers to adopt similar ideas in hallucination detectionmarks2023geometry ; burger2024truth ; How-to-Steer . For instance, HaloScopeHaloScope leverages unlabeled embeddings and applies SVD to identify key subspace directions, followed by probing to link these directions to hallucinations. Yet, probing-based methods often rely on predefined supervised labels, making them less generalizable when feature dimensions are large or category priors are incomplete. Another line of work approaches hallucination detection from the perspective of output consistency. EigenScorechen2024inside quantifies semantic agreement through covariance eigenvalues, while Farquhar et al.farquhar2024detecting utilize clustering and semantic entropy to detect hallucinations. These methods are effective in practice but may suffer from misclassification due to their inability to exploit internal reasoning information.
@@ 3.2 Direct Sum Decomposition of Hidden State Space
To empirically validate the existence and functional role of the reasoning subspace, we design a Reasoning Patch experiment in Appendix E. This experiment demonstrates that the reasoning subspace $\mathcal{S}_{Reasoning}$ indeed captures critical intermediate reasoning information by showing that patching reasoning components from correct solutions can effectively rectify erroneous reasoning trajectories while preserving semantic coherence.
@@ 4.3 Construction of Semantic and Reasoning Subspaces via Low-Rank Approximation
Specifically, based on Equation 7, for any $k<rank(W_{unemb})$, the Eckart–Young–Mirsky theorem Eckart_Young_1936 ; greenacre2022principal gives the best rank-$k$ approximation $W_{k}$ of $W_{unemb}$ under the Frobenius norm as:
@@ 4.4 Hallucination Detection via Reasoning subspace Projection
where $\theta$ denotes the parameters of the hallucination detector. ${{g}_{\theta}}\left({proj}_{R}\left(h_{l}^{(i)}\right)\right)$ represents the hallucination score of the $i$-th token, and $g_{\theta}\left(y|x\right)\in\left[0,1\right]$ is the score for the entire QA pair. We optimize the detector using the Binary Cross-Entropy Loss Goodfellow-et-al-2016 :
@@ 5.1 Experimental Setup
Datasets and models. Our experiments cover four generative question answering (QA) tasks, including three open-domain dialogue QA datasets—NQ OpenNQ_Open , TruthfulQATruthfulQA (generation task), and TriviaQATriviaQA —and one reading comprehension dataset, TyDiQA-GP (English)TyDiQA . To assess the effectiveness and generality of our proposed framework, we conduct evaluations using two widely adopted open-source foundation models: Qwen-2.5-7B-InstructQwen2.5 and LLaMA-3.1-8BLlama_3 . More dataset and inference details are provided in Appendix A.
@@
Baseline Methods. HARP is compared with several mainstream hallucination detection methods, including PerplexityPerplexity , LN-EntropyLN_Entropy , Semantic Entropyfarquhar2024detecting , Lexical SimilarityLexical_Similarity , EigenScorechen2024inside , and HaloScopeHaloScope .
Correctness Measurement. Following Chen et al.chen2024inside , correctness is determined based on ROUGE-L and semantic similarity between generated and reference answers. Semantic similarity is computed using the BLEURT modelBLEURT ; How-to-Steer . An answer is considered correct if its ROUGE-L score exceeds 0.7 or its semantic similarity exceeds 0.5.
@@
LN-Entropy
77.7 63.6 80.2 47.1
Semantic Entropy
77.7 60.0 76.1 68.6
Lexical Similarity
77.8 63.9 76.9 60.3
EigenScore
78.9 63.8 76.2 74.8
@@
LN-Entropy
52.7 62.5 55.8 48.8
Semantic Entropy
60.7 59.4 68.7 62.2
Lexical Similarity
60.9 49.1 71.0 69.5
EigenScore
56.7 45.3 69.1 82.4
@@ 5.2 Main Results
Table 1 summarizes the AUROC scores (in %) of various hallucination detection methods across four QA datasets, using Qwen-2.5-7B-Instruct and LLaMA-3.1-8B as backbone models. Several key findings emerge from these results. (1) HARP consistently outperforms all baseline methods across all datasets and models, often by a significant margin. For instance, on TriviaQA, HARP achieves AUROC scores of 92.8% on Qwen and 92.9% on LLaMA, yielding improvements of +7.5% and +16.6%, respectively, over the second-best method, demonstrating its robustness and scalability across architectures and data characteristics. (2) Baseline methods such as Perplexity and HaloScope perform competitively on simpler datasets like TriviaQA, where answers are often limited to one or two tokens, but their performance deteriorates sharply on more complex datasets such as TyDiQA, which contains long contexts and accompanying documents. In contrast, HARP maintains high AUROC scores of 88.4% on Qwen and 86.6% on LLaMA in these challenging settings, highlighting its ability to handle reasoning-intensive and context-rich inputs. (3) Sampling-based methods, such as Semantic Entropy, Lexical Similarity, and EigenScore, incur higher computational costs but still fail to achieve comparable performance, whereas HARP’s single-pass approach provides both superior efficiency and accuracy. In addition, Table 2 reports the number of known and unknown questions for Qwen-2.5-7B-Instruct across the four datasets, reflecting the model’s varying answering capabilities on these benchmarks. Collectively, these findings validate the effectiveness, robustness, and practical utility of HARP for hallucination detection in diverse QA scenarios.
Table 2: Distribution of known and unknown questions across four QA datasets. A question is classified as Known if the model’s knowledge state contains the correct answer according to the criterion in Equation 1, and as Unknown if none of the 10 candidate responses contain the correct answer.
Dataset Known Unknown
TruthfulQA 636 181
TyDiQA 402 38
TriviaQA 6225 3735
NQ-open 293 3317
@@ 5.3 More Analysis
Selection of Hallucination Score Threshold. In practice, it is necessary to set a hallucination score threshold $\alpha$ so that $\widehat{G}(y|x)$ produces a clear binary decision. As shown in 6(a) and 6(b), when $\alpha$ is between 0.2 and 0.8, both detection accuracy and F1 score remain high, indicating a substantial separation between normal and hallucinated answers under $\widehat{G}$. To align with common expectations for a binary classifier, we set $\alpha=0.5$, where $\widehat{G}(y|x)=\mathbb{I}\left[g_{\theta}(y|x)>0.5\right]$.
Figure 6: (a) Effect of hallucination score threshold on detection accuracy. (b) Effect of hallucination score threshold on detection F1 score.
Robustness Experiments. To apply HARP in real-world scenarios, we examined its performance under distribution shifts between training and test sets. We trained the hallucination detector on a source dataset $s$ and evaluated it on different target datasets $t$. Figure 7 shows that HARP generalizes well across multiple target datasets. Notably, when trained on TriviaQA, its accuracy on NQ Open is nearly identical to directly training on NQ Open, demonstrating HARP’s strong robustness and cross-distribution adaptability.
Figure 7: Cross-dataset generalization. “(s)” indicates the source dataset used for training the hallucination detector; “(t)” indicates the target dataset.
@@ 6 Conclusion
In this study, we introduced HARP, a novel hallucination detection method that leverages only reasoning information as input features, achieving high detection accuracy while maintaining strong robustness. First, we showed that the hidden state space admits a direct-sum decomposition into a semantic subspace and a reasoning subspace, and that the Unembedding layer can effectively separate these two components. Building on this, we applied singular value decomposition to the parameters of the Unembedding layer and, following the Eckart–Young–Mirsky theorem, approximated $W_{unemb}$ with its best rank-$k$ approximation $W_{k}$. Setting $k=d\times 95\%$, we identified basis vectors for both the semantic and reasoning subspaces that align with empirical observations. Furthermore, we empirically validated that the reasoning subspace effectively captures intermediate reasoning information through the Reasoning Patch experiment detailed in Appendix E. Finally, HARP constructs an accurate and efficient hallucination detector by using the projections of hidden states in the reasoning subspace as input features. Experiments show that HARP significantly outperforms existing mainstream hallucination detection methods and maintains robustness under distribution shifts across datasets. In addition, we present a proof-of-concept demonstration of hallucination mitigation using our framework in Appendix D and aim to inspire future research in this direction.
@@ Appendix D Mitigating Hallucinations by Modifying Reasoning Subspace Components
Appendix E Verification of Reasoning Information in the Reasoning Subspace
To verify that the components of hidden states lying in the reasoning subspace indeed encode internal reasoning information, we design a controlled experiment consisting of three input conditions (Figure 9). These conditions isolate the effect of the reasoning subspace while keeping all other factors unchanged.
Figure 9: Experimental design illustrating the three conditions used to verify the causal role of the reasoning subspace. (A)Normal: the model receives only the question and produces an incorrect answer. (B)CoT: a chain-of-thought is prepended, enabling multi-step reasoning and a correct answer. (C)Reasoning Patch: no CoT is provided, but the reasoning-subspace components of hidden states at all layers are replaced with those from the CoT run, causing the model to generate reasoning steps and arrive at the correct answer.
(A) Normal: direct question input.
In the first condition, we feed the model only the question without any chain-of-thought (CoT) guidance. The model typically produces an incorrect answer. Let the hidden state be
$h_{\cdot}^{Normal}=h_{\cdot,Semantic}^{Normal}\;+\;h_{\cdot,Reasoning}^{Normal}.$
(B) CoT: prepend chain-of-thought.
In the second condition, we prepend a chain-of-thought[32] to the input. The model now first generates intermediate reasoning steps and then outputs the correct answer. The hidden state is
$h_{\cdot}^{CoT}=h_{\cdot,Semantic}^{CoT}\;+\;h_{\cdot,Reasoning}^{CoT}.$
(C) Reasoning Patch: replace reasoning components at all relevant layers.
The third condition serves as the key causal intervention. The input text is identical to condition (A); however, at every decoder layer that contributes to the representation of a token, we replace only the reasoning-subspace component of the hidden state with the corresponding component extracted from condition (B). Formally, for all layers along the forward-pass trajectory of token $t$, we apply:
$h^{Patch}_{\cdot}=h^{Normal}_{\cdot,Semantic}\;+\;h^{CoT}_{\cdot,Reasoning}.$
Thus, semantic information is preserved at every layer, while the reasoning components across all intermediate layers are substituted with those from the CoT run. This ensures that the patched forward pass follows the CoT reasoning trajectory throughout the entire decoder stack.
Key result.
We evaluate the effectiveness of the proposed Reasoning Patch on mathematical reasoning benchmarks such as GSM8K[33], using both few-shot CoT and zero-shot CoT to extract the reasoning-subspace components.
Figure 10: Reasoning Patch experiment using few-shot chain-of-thought supervision.
Figure 10 presents the Qwen2.5-7B-Instruct outputs under the three conditions (A)–(C) when the reasoning components of condition (C) are derived from few-shot CoT, with the full prompts shown in Table 5. We observe that, even though condition (C) receives no CoT text in the input, injecting the CoT-derived reasoning-subspace components reliably triggers the model to follow a “reason-then-answer” generation pattern. As a result, the model transitions from an incorrect answer in (A) to a correct, multi-step reasoning process in (C), demonstrating that the patched reasoning trajectory causally determines the emergence of correct step-by-step reasoning.
Table 5: few-shot chain-of-thought prompt.
Figure 11 shows the corresponding results when the reasoning components are extracted from zero-shot CoT. Remarkably, even though condition (C) does not contain the zero-shot instruction (e.g., “Answer the following question step by step to the best of your ability.”), the patched model nonetheless produces a coherent step-by-step reasoning chain before giving the final answer. Interestingly, in this setting the original CoT run in condition (B) makes an arithmetic mistake and outputs an incorrect final answer; however, condition (C)—which inherits only the reasoning-subspace components rather than the explicit token sequence—does not reproduce this error and instead produces the correct result. This highlights that the reasoning subspace captures the structural reasoning trajectory without being constrained by the semantic information in the CoT prompt.
Figure 11: Reasoning Patch experiment using zero-shot chain-of-thought prompting.
Together, these results provide compelling evidence that the reasoning subspace encodes causally meaningful internal reasoning information, and that injecting its components is sufficient to induce coherent multi-step reasoning even in the absence of explicit CoT prompting.
Appendix F Computational Complexity of SVD
To construct the reasoning subspace, we perform singular value decomposition (SVD) on a matrix $M\in\mathbb{R}^{n\times d}$, where $n$ denotes the vocabulary size and $d$ is the dimensionality of the hidden representation. In typical large language models, the matrix is tall and skinny with $n\gg d$ (e.g., for Qwen2.5-7B, n=152,064n=152{,}064 and d=3,584d=3{,}584). The computational complexity of SVD depends on these matrix dimensions as well as whether a full or truncated decomposition is applied.
Time Complexity.
For a full SVD on an $n\times d$ matrix, the time complexity is
$O\!\left(\min(nd^{2},\,n^{2}d)\right).$
Since the vocabulary size is typically much larger than the hidden dimension, the dominant term becomes
$O(nd^{2}),$
which makes full SVD computationally expensive in practice. For truncated SVD that retains only the top-$k$ singular directions, the complexity reduces to
$O(ndk),$
particularly when using iterative or randomized SVD algorithms. Such approximations are crucial for scaling to vocabularies of realistic size.
Space Complexity.
Storing the matrix $M$ requires
$O(n^{2}+nd+d^{2})$
memory. The truncated singular vectors $U\in\mathbb{R}^{n\times k}$ and $V\in\mathbb{R}^{d\times k}$ introduce an additional
$O((n+d)k)$
space overhead. Because $n$ is very large in modern LLMs, the memory is dominated by storing $U$.
SVD Resource Consumption.
To quantify the practical resource requirements of performing SVD on the unembedding layer, Table 6 summarizes the wall-clock time, the peak memory consumption during the SVD computation, and the additional memory introduced by truncated SVD across several representative models. The evaluation covers models of different scales—including Qwen2.5-7B-Instruct, LLaMA-3.1-8B, Qwen2.5-72B-Instruct, and the MoE model Qwen3-235B-A22B-Instruct-2507-FP8 [34]—using an H100 80GB GPU.
Table 6: SVD computation cost on the unembedding layer using an H100 80GB GPU. We report wall-clock time, memory required during the SVD computation, and the additional memory by SVD.
Model Unembedding Shape Time Peak Memory Extra Memory
Qwen2.5-7B-Instruct $152{,}064\times 3{,}584$ 1.30s 8.37GB 0.02GB
LLaMA-3.1-8B $128{,}256\times 4{,}096$ 1.60s 8.08GB 0.03GB
Qwen2.5-72B-Instruct $152{,}064\times 8{,}192$ 9.83s 19.22GB 0.13GB
Qwen3-235B-A22B-Instruct-2507-FP8 (MoE) $151{,}936\times 4{,}096$ 1.70s 9.55GB 0.03GB
Singular Value Distribution in Larger Models.
Figure 12 shows the singular value distribution of the unembedding layers for larger models, including Qwen2.5-72B-Instruct and Qwen3-235B-A22B-Instruct-2507-FP8 (MoE). The trend of singular value decay is consistent with that observed for Qwen2.5-7B-Instruct and LLaMA-3.1-8B (4(a)), indicating that our method can be directly applied to larger models.
Figure 12: Singular value distributions of Wunemb after SVD, with hidden state dimensions of 8192 for Qwen2.5-72B-Instruct and 4096 for Qwen3-235B-A22B-Instruct-2507-FP8 (MoE).